OpenStudy (safyousuff97):

Jake is comparing the prices of two laundry service companies. Company A charges $20 per load and an additional $8 as service charges. Company B charges $24 per load and an additional $4 as service charges. Part A: Write equations to represent Company A's and Company B's total laundry charges for a certain number of loads. Define the variable used in the equations. Part B: Which company would charge less for 5 loads of laundry? Justify your answer. Part C: How much money is saved by using the services of Company A instead of Company B to clean 9 loads of laundry?

OpenStudy (safyousuff97):

@TheSmartOne

TheSmartOne (thesmartone):

So do you know how to write an equation, yousuf?

OpenStudy (safyousuff97):

i'm not very good at writing equations but I will try

TheSmartOne (thesmartone):

So do you know what the constant could be for the equations?

OpenStudy (safyousuff97):

would it be 20 and 8

TheSmartOne (thesmartone):

No, those numbers are for PER LOAD The other two numbers are a one time fee for any amount of load Do you understand? :)

OpenStudy (safyousuff97):

oh ok

OpenStudy (safyousuff97):

i'm still confused though

TheSmartOne (thesmartone):

So, for part a we can make two equations Company A's equation is y = 20x + 8 Company B's equation is y = 24x + 4 Do you agree? :)

OpenStudy (safyousuff97):

yes

TheSmartOne (thesmartone):

Now for part b, do you know what to do?

OpenStudy (safyousuff97):

i do the equation now right

TheSmartOne (thesmartone):

you plug in 5 into both equations and see which company has a smaller amount for y :)

TheSmartOne (thesmartone):

do it yourself, but post your work so I can make sure I'll wait for you reply :)

OpenStudy (safyousuff97):

ok i will

TheSmartOne (thesmartone):

:)

OpenStudy (safyousuff97):

im stuck on how place the 5 in the equation

OpenStudy (safyousuff97):

y=20x+8 5 -5 am i doing it wrong?

TheSmartOne (thesmartone):

yea

TheSmartOne (thesmartone):

you replace x with 5

TheSmartOne (thesmartone):

so y = 20(5) + 8 y = 100 + 8 y = ?

OpenStudy (safyousuff97):

ok i'm sorry but i'm really dumb and i have know idea what to do but is y-=108

OpenStudy (safyousuff97):

y=108

TheSmartOne (thesmartone):

yes

TheSmartOne (thesmartone):

now do it for the other equation :)

OpenStudy (safyousuff97):

ok

OpenStudy (safyousuff97):

is it 95

TheSmartOne (thesmartone):

how did you get that?

OpenStudy (safyousuff97):

by doing this y=20(5)+8 5 -100 95

OpenStudy (anonymous):

do you still need help with this?

OpenStudy (safyousuff97):

i don't know what to do, i'm very confused

OpenStudy (safyousuff97):

i don't know the steps for this problem

TheSmartOne (thesmartone):

try this y = 24(5) + 4 solve for y :)

OpenStudy (safyousuff97):

is it 124

OpenStudy (anonymous):

hold on I will help you with all parts k?

OpenStudy (safyousuff97):

ok thx

TheSmartOne (thesmartone):

why?

TheSmartOne (thesmartone):

when I'm already helping..

TheSmartOne (thesmartone):

and that's correct @safyousuff97

OpenStudy (safyousuff97):

ok thats for part a right

TheSmartOne (thesmartone):

so which company is cheaper for 5 loads? Company A charges 108 for 5 loads Company B charges 124 for 5 loads

TheSmartOne (thesmartone):

this is for part b, we completed part a already :)

OpenStudy (safyousuff97):

company a

TheSmartOne (thesmartone):

correct that's your answer for part b :)

OpenStudy (anonymous):

Company A $20 per load $8 service Company B $24 per load $4 service PART A: Let l be the number of loads Company A = 20l + 8 Company B = 24l + 4 PART B: Company A = 20*5 + 8 = $108 Company B = 24*5 + 4 = $124 Company A charges less PART C: Company A = 20*9 + 8 = $188 Company B = 24*9 + 4 = $220 220 - 188 = 32 You save $32.

TheSmartOne (thesmartone):

and now do what we did for part b for part c but instead of using 5, use 9

TheSmartOne (thesmartone):

gee, you really took the learning out of the whole question @*Nikki*

OpenStudy (anonymous):

lol

OpenStudy (safyousuff97):

@*Nikki* thank you but

OpenStudy (safyousuff97):

I wanted you to explain how to do the problem

OpenStudy (safyousuff97):

thank you anyway

OpenStudy (safyousuff97):

how do i give 2 medals

OpenStudy (anonymous):

thats not possible unless you undo one!! :D

TheSmartOne (thesmartone):

it is possible through a glitch, but that would be exploiting it :)

OpenStudy (safyousuff97):

oh ok i'm sorry @*Nikki* i'm won't be able to give a medal to you

OpenStudy (anonymous):

well then next time u can!! =]

OpenStudy (safyousuff97):

ok ,thank you for your help :D

TheSmartOne (thesmartone):

lol

OpenStudy (safyousuff97):

thank you @TheSmartOne for explaining the question

TheSmartOne (thesmartone):

anytime :D

OpenStudy (safyousuff97):

k bye :D